OpenStudy (anonymous):

Do the sides of root7, 3, 4 form a right triangle?

OpenStudy (anonymous):

Check to see if it fits in the Pythagorean Theorem for right triangles: \[a^2+b^2=c^2\], when a=7,b=3, and 4=c.

OpenStudy (anonymous):

We come out with \[\sqrt{7}^2+3^3=4^2\]

OpenStudy (anonymous):

Which is correct. So it is a right triangle.

OpenStudy (anonymous):

that actually helped alot thank you :)